# On the dynamics of leapfrogging vortex rings

### v1

@@ 5.2. Limit trajectories
First, we consider $\widetilde{Y}_{i}$ defined in (1.10). As it is a system independent of $\varepsilon$, it is clear by continuity of the trajectories that there exists $T_{X}\in(0,1]$ such that
@@
Next, we remark that $Y_{i}^{\varepsilon}$ is given by
@@
The map $\mathcal{F}^{\varepsilon}_{1}$ is Lipschitz continuous when $\mathcal{X}^{\varepsilon},\widetilde{\mathcal{X}}^{\varepsilon}\in\prod_{i}\mathcal{A}_{\varepsilon}^{X^{\varepsilon}_{i,0,r}}\left(d_{0},h_{0}\right)$ with constant $K=\mathcal{O}(1)$, so by Gronwall lemma (see for instance the variant in Lemma A.2), we get that for every $t\leqslant T_{\varepsilon}$,
$\left|\mathcal{Y}^{\varepsilon}(t)-\widetilde{\mathcal{Y}}^{\varepsilon}(t)\right|\leqslant\left(\frac{CT_{X}}{\sqrt{|\ln\varepsilon|}}+0\right)e^{KT_{X}},$
@@
Provided $T_{0}\leqslant\min(1/(26C^{\prime}_{F}),T_{X})$, there exist $\varepsilon_{0}>0$ and $C>0$ independent of $T_{0}$, such that for every $\varepsilon\in(0,\varepsilon_{0}]$, $i\in\{1,\ldots,N\}$ and $t\in[0,T_{\varepsilon}]$, we have that
@@
From this result, we are now in position to prove Theorem 1.1-$(ii)$ on the time interval $t\in[0,T_{\varepsilon}]$.
@@
Provided $T_{0}\leqslant\min(1/(26C^{\prime}_{F}),T_{X})$, there exist $\varepsilon_{0}>0$ and $C_{0}>0$ independent of $T_{0}$, such that for every $\varepsilon\in(0,\varepsilon_{0}]$, $i\in\{1,\ldots,N\}$, we have that
@@
By Proposition 5.4 we have that
@@ 6. Strong localization
All the above results are valid up to $T_{\varepsilon}$, defined in (3.6), in particular Part (ii) of Theorem 1.1. The goal of this section is to get Part (ii) of Theorem 1.1, but also to state that $T_{\varepsilon}=T_{0}$, namely we will prove the following.
@@
For $h_{0}$ fixed, there exists $T_{0}>0$ and $\varepsilon_{0}>0$ such that for every $\varepsilon\in(0,\varepsilon_{0})$, every $t\leqslant T_{\varepsilon}$ and $i\in\{1,\ldots,N\}$, we have that
@@
We start by establishing the strong localization in the radial direction, which requires sharp estimates and a two-step iteration process as in [11], then turn to the vertical strong localization by a simpler version of the same method.
@@ 6.1. Strong localization in the radial direction
Let $r_{0}^{i,\varepsilon}=X_{i,0,r}^{\varepsilon}$. For any $R>0$, let
$\mathcal{U}^{i,\varepsilon}_{R}=\{(r,z)\in{\mathbb{R}}^{2}_{+},\;|r-r_{0}^{i,\varepsilon}|<R\},$
@@
$m_{t}^{i,\varepsilon}(R)=\int_{(\mathcal{U}^{i,\varepsilon}_{R})^{c}}x_{r}\omega^{\varepsilon}_{i}(t,x)\mathrm{d}x$
and let
$\mu_{t}^{i,\varepsilon}(\eta,R)=\int_{{\mathbb{R}}^{2}}\big{(}1-W_{R,\eta}(x_{r}-r_{0}^{i,\varepsilon})\Big{)}x_{r}\omega^{\varepsilon}_{i}(t,x)\mathrm{d}x$
@@
$\Big{|}x^{i,\varepsilon,0}_{r}-r_{0}^{i,\varepsilon}\Big{|}=R_{t}^{i,\varepsilon}:=\max\Big{\{}\Big{|}x_{r}-r_{0}^{i,\varepsilon}\Big{|}\,,\,x\in\operatorname{supp}\omega^{\varepsilon}_{i}(t)\Big{\}}.$
Let $s\mapsto x^{i,\varepsilon,t}(s)$ be the Lagrangian trajectory passing by $x^{i,\varepsilon,0}$ at time $t$, i.e. the solution of
[…]
With these notations, we have the following.
Lemma 6.2.
There exists $C>0$ independent of $T_{0}$ such that
Proof.
We compute that
and thus, recalling that $x^{\varepsilon,t}(t)=x^{\varepsilon,0}$,
$\leqslant\frac{C}{|\ln\varepsilon|}\int_{{\mathbb{R}}^{2}_{+}}\frac{y_{r}\omega^{\varepsilon}(t,y)}{\big{|}x^{\varepsilon,0}-y\big{|}}\mathrm{d}y+\frac{C}{\sqrt{|\ln\varepsilon|}}.$
We now split the integral on $\mathcal{U}^{\varepsilon}_{R_{t}^{\varepsilon}/2}$ and +$\mathcal{{\mathbb{R}}}^{2}_{+}\setminus\mathcal{U}^{\varepsilon}_{R_{t}^{\varepsilon}/2}.$ If $y\in\mathcal{U}^{\varepsilon}_{R_{t}^{\varepsilon}/2}$, then $|x^{\varepsilon,0}-y|\geqslant R_{t}^{\varepsilon}/2$, and thus
On the complement, we use the following property [^2] See for instance [32] when $\Omega={\mathbb{R}}^{2}$ and $f\in L^{1}\cap L^{\infty}({\mathbb{R}}^{2})$, and next apply this for $f\mathds{1}_{\Omega}$.: for every $\Omega\subset{\mathbb{R}}^{2}$ and $f\in L^{1}\cap L^{\infty}(\Omega)$
where $C$ is independent of $\Omega$, to obtain that
+$\int_{{\mathbb{R}}^{2}_{+}\setminus\mathcal{U}^{\varepsilon}_{R_{t}^{\varepsilon}/2}}\frac{y_{r}\omega^{\varepsilon}(t,y)}{|x^{\varepsilon,0}-y|}\mathrm{d}y\leqslant\frac{C}{\varepsilon}\sqrt{m_{t}^{\varepsilon}(R_{t}^{\varepsilon}/2)}.$
We now work on estimating the quantity $m_{t}^{i,\varepsilon}(R_{t}^{i,\varepsilon}/2)$. To this end we use a double iterative method inspired by [11].
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\mu_{t}^{i,\varepsilon}(R,\eta)\leqslant C\left(\frac{1}{\eta\sqrt{\ln\varepsilon}}+\frac{m_{t}^{i,\varepsilon}(R/2)}{\eta^{2}|\ln\varepsilon|}\right)m_{t}^{i,\varepsilon}(R)$
@@
Recalling that $u_{L}^{\varepsilon}\cdot e_{r}=0$, we compute that
$\frac{\mathrm{d}}{\mathrm{d}t}\mu_{t}^{\varepsilon}(R,\eta)=\frac{-1}{|\ln\varepsilon|}\int W_{R,\eta}^{\prime}(x_{r}-r^{\varepsilon}_{0})e_{r}\cdot(u_{K}^{\varepsilon}+u_{R}^{\varepsilon}+F^{\varepsilon})(t,x)x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x=A_{1}+A_{2}$
@@
We start with $A_{1}$. Recalling that $\|u_{R}^{\varepsilon}\|_{L^{\infty}}\leqslant C$ and $\|F^{\varepsilon}\|_{L^{\infty}}\leqslant C\sqrt{|\ln\varepsilon|}$, and that $W_{R,\eta}^{\prime}=0$ out of $[R,R+\eta]$ we have that
@@
We now compute $A_{2}$. First, we expend $u_{K}^{\varepsilon}$ and get that
@@
then we symmetrize to get that
@@
$|A_{22}+A_{23}|\leqslant\frac{C}{\eta^{2}|\ln\varepsilon|}\iint_{A_{R}^{c}\times A_{R/2}^{c}}x_{r}\omega^{\varepsilon}(t,x)y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}x\mathrm{d}y\leqslant\frac{C}{\eta^{2}|\ln\varepsilon|}m_{t}^{\varepsilon}(R)m_{t}^{\varepsilon}(R/2),$
@@
by choosing $T_{0}$ small enough. From Lemma 5.3 and up to choose $\varepsilon$ small enough, we have
@@
provided $R\in\left[\frac{d_{0}}{128\sqrt{|\ln\varepsilon|}},\frac{d_{0}}{64\sqrt{|\ln\varepsilon|}}\right]$ and $\varepsilon_{0}$ is small enough. Corollary 5.5 then implies that
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\mu_{t}^{\varepsilon}(R,\eta)\leqslant C\left(\frac{1}{\eta\sqrt{\ln\varepsilon}}+\frac{\ln|\ln\varepsilon|}{\eta^{2}|\ln\varepsilon|^{2}}\right)m_{t}^{\varepsilon}(R),$
for any $R\in\left[\frac{d_{0}}{128\sqrt{|\ln\varepsilon|}},\frac{d_{0}}{64\sqrt{|\ln\varepsilon|}}\right]$. Now we set $R_{0}=\frac{d_{0}}{64\sqrt{|\ln\varepsilon|}}$, $n=[\ln|\ln\varepsilon|]-1$ and $\eta=\frac{R_{0}}{2n}$ and we obtain that
@@
$\mu_{t}^{\varepsilon}(R,\eta)\leqslant\mu_{0}(R,\eta)+C\,n\int_{0}^{t}m_{s}(R)\mathrm{d}s$
@@
$\mu_{t}^{\varepsilon}(R,\eta)\leqslant\mu_{0}(R,\eta)+Cn\int_{0}^{t}\mu_{s}(R-\eta,\eta)\mathrm{d}s,$
@@
$\mu_{t}^{\varepsilon}(R_{0},\eta)\leqslant\mu_{0}^{\varepsilon}(R_{0},\eta)+Cnt\mu_{0}^{\varepsilon}(R_{0}-\eta,\eta)+(Cn)^{2}\int_{0}^{t}\int_{0}^{s_{1}}\mu_{s_{2}}^{\varepsilon}(R_{0}-2\eta,\eta)\mathrm{d}s_{2}\mathrm{d}s_{1}$
and iterating $n$ times eventually leads to
$\mu_{t}^{\varepsilon}(R_{0},\eta)\leqslant\sum_{j=0}^{n-1}\mu_{0}^{\varepsilon}(R_{0}-j\eta,\eta)(Cnt)^{j}+(Cn)^{n}\int_{0}^{t}\ldots\int_{0}^{s_{n-1}}\mu_{s_{n}}^{\varepsilon}(R_{0}/2,\eta)\mathrm{d}s_{n}\ldots\mathrm{d}s_{1}$
and therefore, since $\mu_{0}(0/2)=0$ and $\mu_{s}^{\varepsilon}\leqslant\gamma$ for every $s$, we obtain that
$\mu_{t}^{\varepsilon}(R_{0},\eta)\leqslant\frac{(Cnt)^{n}}{n!}.$
@@
$|\ln\varepsilon|^{\ell}\mu_{t}^{\varepsilon}(R_{0},\eta)$ $\leqslant\frac{1}{\sqrt{\pi n}}\exp\Big{(}n\big{(}\ln(Cnt)-\ln(n)+1\big{)}+\ell\ln|\ln\varepsilon|\Big{)}$
$\leqslant\frac{2}{\sqrt{\pi\ln|\ln\varepsilon|}}\exp\Big{(}\ln|\ln\varepsilon|\big{(}\ln(Ct)+1+\ell\big{)}\Big{)}.$
We see that if $T_{0,\ell}\leqslant\frac{1}{Ce^{\ell+1}}$, then this expression goes to 0 as $\varepsilon\to 0$ uniformly in $t\in[0,T_{\varepsilon}]$. ∎
@@
By taking this time $R_{0}=\frac{d_{0}}{16\sqrt{|\ln\varepsilon|}}$, $\eta=\frac{R}{2n}$ and $n=[|\ln\varepsilon|]-1$, we get again that
@@
$\mu_{t}^{\varepsilon}(R_{0},\eta)\leqslant\frac{(Cnt)^{n}}{n!}.$
@@
$\varepsilon^{-\ell}\mu_{t}^{\varepsilon}(R_{0},\eta)$ $\leqslant\frac{1}{\sqrt{\pi n}}\exp\Big{(}n\big{(}\ln(Cnt)-\ln(n)+1\big{)}+\ell|\ln\varepsilon|\Big{)}$
$\leqslant\frac{2}{\sqrt{\pi|\ln\varepsilon|}}\exp\Big{(}|\ln\varepsilon|\big{(}\ln(Ct)+1+\ell\big{)}\Big{)}.$
which again goes to 0 if $T_{0,\ell}\leqslant\frac{1}{Ce^{\ell+1}}$. ∎
We are now able to conclude the radial part of Proposition 6.1.
@@
Indeed, if it isn’t the case, by letting $t_{2}\in[t_{0},t_{1}]$ be the first time when $R_{t_{2}}^{\varepsilon}=f(t_{2})$, then there exists a point $x^{\varepsilon,0}\in\operatorname{supp}\omega^{\varepsilon}(t_{2})$ such that $|x^{\varepsilon,0}-r^{\varepsilon}_{0}|=R_{t_{2}}^{\varepsilon}$. However Lemma 6.2 is then in contradiction with the fact that $t_{2}$ is the first such time, due to the fact that $f(t_{0})>R_{t_{0}}^{\varepsilon}$ and

### v2

@@ 5.2. Limiting trajectories
First, we consider $\widetilde{Y}_{i}$ defined in (1.8). Given the independence of $\varepsilon$ of the respective system, it is clear by continuity of the trajectories that there exists $T_{X}\in(0,1]$ such that
@@
Next, we remark that by definition $Y_{i}^{\varepsilon}$ is given by
@@
The map $\mathcal{F}^{\varepsilon}_{1}$ is Lipschitz continuous when $\mathcal{X}^{\varepsilon},\widetilde{\mathcal{X}}^{\varepsilon}\in\prod_{i}\mathcal{A}_{\varepsilon}^{X^{\varepsilon}_{i,0,r}}\left(d_{0},h_{0}\right)$ with constant $K=\mathcal{O}(1)$, so by Gronwall lemma (see for instance the variant in Lemma A.2), we get that for every $t\leqslant\widehat{T}_{\varepsilon}$,
$\left|\mathcal{Y}^{\varepsilon}(t)-\widetilde{\mathcal{Y}}(t)\right|\leqslant\left(\frac{CT_{X}}{\sqrt{|\ln\varepsilon|}}+0\right)e^{KT_{X}},$
@@
Provided $T_{0}\leqslant\min(1/(24C^{\prime}_{F}),T_{X})$, there exist $\varepsilon_{0}>0$ and $C>0$ independent of $T_{0}$, such that for every $\varepsilon\in(0,\varepsilon_{0}]$, $i\in\{1,\ldots,N\}$ and $t\in[0,T_{\varepsilon}]$, we have that
@@
From this result, we are now in position to prove Theorem 1.1-$(i)$ on the time interval $t\in[0,T_{\varepsilon}]$.
@@
Provided $T_{0}\leqslant\min(1/(24C^{\prime}_{F}),T_{X})$, there exist $\varepsilon_{0}>0$ and $C_{0}>0$ independent of $T_{0}$, such that for every $\varepsilon\in(0,\varepsilon_{0}]$, $i\in\{1,\ldots,N\}$, we have that
@@
By Proposition 5.4 and as $h_{0}\leqslant\frac{1}{2}$ (see Remark 4.12), we have that
@@ 6. Strong localization
All the above results are valid up to $T_{\varepsilon}$, defined in (3.6), in particular Part (i) of Theorem 1.1. The goal of this section is to show Part (ii) of Theorem 1.1, but also to state that $T_{\varepsilon}=T_{0}$, namely we will prove the following.
@@
For $h_{0}$ fixed, there exists $T_{0}>0$ and $\varepsilon_{0}>0$ such that for every $\varepsilon\in(0,\varepsilon_{0}]$, every $t\leqslant T_{\varepsilon}$ and $i\in\{1,\ldots,N\}$, we have that
@@
We start by establishing the strong localization in the radial direction, which requires precise estimates and a two-step iteration process inspired by [11], then turn to the vertical strong localization proven by a simpler version of the same method.
This strong localization method was introduced by Marchioro and Pulvirenti in the nineties, first published in [44], see also [45], in the context of $2D$-Euler flows. It has since proven to be a powerful tool for the study of concentrated vorticities and their stability and been deployed in many important articles. The approach has mainly been developed by Marchioro and his collaborators including adaptations to the $3D$ setting of vortex rings in various regimes. The arguments have been gradually refined and substantially improved in a long series of papers and noways reached an apparent level of complexity that may make it seem difficult to understand. Our aim in this section is to present this powerful approach in the most pedagogical and self-contained way possible, allowing the reader who has skimmed the previous sections to learn this tool.
@@ 6.1. Strong localization in the radial direction
To shorten the notation, let us denote for this section $r_{0}^{i,\varepsilon}=X_{i,0,r}^{\varepsilon}$. Strong localization means controlling the size of the support of the vorticity, i.e. of the following quantity
$R_{t}^{i,\varepsilon}:=\max\Big{\{}\Big{|}x_{r}-r_{0}^{i,\varepsilon}\Big{|}\,,\,x\in\operatorname{supp}\omega^{\varepsilon}_{i}(t,\cdot)\Big{\}}.$
Although the natural definition would be to use the Euclidian distance in ${\mathbb{R}}^{2}$, the presence of the local induction term $v_{L}^{\varepsilon}$ in 3D axisymmetric flows, as noticed in [8], makes this problem anisotropic and forces us to treat independently the radial and the vertical directions. Because of possible filamentation, we obtain a stronger control in the radial direction rather than in the vertical direction. Let us mention that we could also define this radius by measuring the distance to the center of mass $b^{\varepsilon}_{i}(t)$ rather than to $X_{i,0,r}$. However, in the current setting, it would not improve the obtained estimates but would result in longer computations.
To estimate the evolution of $R_{t}^{i,\varepsilon}$, we now fix a time $t\leqslant T_{\varepsilon}$ and an index $i\in\{1,\ldots,N\}$, and focus on a fluid particle trajectory from the support of $\omega_{i}^{\varepsilon}$ that reaches at time $t$ the farthest distance from the initial position (in the radial direction). Namely, we consider $x^{i,\varepsilon,0}\in{\mathbb{R}}^{2}_{+}$ such that
$\Big{|}x^{i,\varepsilon,0}_{r}-r_{0}^{i,\varepsilon}\Big{|}=R_{t}^{i,\varepsilon},$
and $s\mapsto x^{i,\varepsilon,t}(s)$ be the Lagrangian trajectory passing by $x^{i,\varepsilon,0}$ at time $t$, i.e., by the transport equation (3.2), the solution of
[…]
The general idea of the proof is to control the velocity of this fluid particle in the radial direction (and hence the growth of $R_{t}^{i,\varepsilon}$) by controlling the mass of the vorticity included in a region located between $R_{t}^{i,\varepsilon}/2$ and $R_{t}^{i,\varepsilon}$. Indeed, the fluid particle we focus on is relatively far from the vortex core, which gives good control over the velocity induced by the latter on the former. Therefore, we prove that estimating the vorticity mass far from the vortex core (hence potentially close to the fluid particle) induces control over its outward velocity. This is the content of the following lemma. Let us introduce for any $R>0$,
$\mathcal{U}^{i,\varepsilon}_{R}:=\{(z,r)\in{\mathbb{R}}^{2}_{+},\;|r-r_{0}^{i,\varepsilon}|<R\},$
@@
$m_{t}^{i,\varepsilon}(R):=\int_{(\mathcal{U}^{i,\varepsilon}_{R})^{c}}x_{r}\omega^{\varepsilon}_{i}(t,x)\mathrm{d}x.$
Lemma 6.2.
There exists $C>0$ independent of $T_{0}$ such that
Proof.
As often in the previous sections, we drop during the proofs the index $i$ in the notation of $\omega_{i}^{\varepsilon}$ and $\Omega_{i}^{\varepsilon}$ since we perform the proofs for a fixed $i\in\{0,\ldots,N\}$ and we derive uniform estimates with respect to this parameter.
We compute that
and thus, recalling that $x^{\varepsilon,t}(t)=x^{\varepsilon,0}$,
$\leqslant\frac{C}{|\ln\varepsilon|}\int_{{\mathbb{R}}^{2}_{+}}\frac{y_{r}\omega^{\varepsilon}(t,y)}{\big{|}x^{\varepsilon,0}-y\big{|}}\mathrm{d}y+\frac{C}{\sqrt{|\ln\varepsilon|}},$
where we have used, as often in Section 4, the decomposition of $u^{\varepsilon}$ (see Lemma 2.6), that $u_{L}^{\varepsilon}\cdot e_{r}=0$, the estimate on $F^{\varepsilon}$ (see Lemma 3.1) and that $r^{*}/2\leqslant x^{\varepsilon,0},y\leqslant 2r^{*}$ (see (3.7)). We now split the integral on $\mathcal{U}^{\varepsilon}_{R_{t}^{\varepsilon}/2}$ and +$\mathcal{{\mathbb{R}}}^{2}_{+}\setminus\mathcal{U}^{\varepsilon}_{R_{t}^{\varepsilon}/2}.$ If $y\in\mathcal{U}^{\varepsilon}_{R_{t}^{\varepsilon}/2}$, then $|x^{\varepsilon,0}-y|\geqslant R_{t}^{\varepsilon}/2$, and thus, by the conservation of the total mass of vorticity (3.3) and as the vorticity is non-negative, we have
On the complement, we use the following property [^3] See for instance [36, Lemma 2.1] when $\Omega={\mathbb{R}}^{2}$ and $f\in L^{1}\cap L^{\infty}({\mathbb{R}}^{2})$, and next apply this for $f\mathds{1}_{\Omega}$.: for every $\Omega\subset{\mathbb{R}}^{2}$ and $f\in L^{1}\cap L^{\infty}(\Omega)$
where $C$ is independent of $\Omega$, to obtain that
+$\int_{{\mathbb{R}}^{2}_{+}\setminus\mathcal{U}^{\varepsilon}_{R_{t}^{\varepsilon}/2}}\frac{y_{r}\omega^{\varepsilon}(t,y)}{|x^{\varepsilon,0}-y|}\mathrm{d}y\leqslant\frac{C}{\varepsilon}\sqrt{m_{t}^{\varepsilon}(R_{t}^{\varepsilon}/2)},$
because the transported initial vorticity is less than $C/\varepsilon^{2}$, see (1.5). ∎
It appears now clear that we need to estimate the mass of the vorticity further than $R^{i,\varepsilon}_{t}/2$, namely we are looking for a weak localization estimate. Unfortunately, Corollary 4.3 derived in Section 4 is not enough because $\ln|\ln\varepsilon|/(\varepsilon^{2}|\ln\varepsilon|)\gg|\ln\varepsilon|$ so we will prove a better estimate for $m_{t}^{i,\varepsilon}(R_{t}^{i,\varepsilon}/2)$.
As we will need to differentiate with respect to $t$, which is difficult because the domain of integration in the definition of $m_{t}^{i,\varepsilon}$ depends on $t$, we introduce a regularization of $m_{t}^{i,\varepsilon}$: let
$\mu_{t}^{i,\varepsilon}(R,\eta):=\int_{{\mathbb{R}}^{2}}\big{(}1-W_{R,\eta}(x_{r}-r_{0}^{i,\varepsilon})\Big{)}x_{r}\omega^{\varepsilon}_{i}(t,x)\mathrm{d}x$
@@
We estimate in the next lemma the time derivative.
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\mu_{t}^{i,\varepsilon}(R,\eta)\leqslant C\left(\frac{1}{\eta\sqrt{|\ln\varepsilon|}}+\frac{m_{t}^{i,\varepsilon}(R/2)}{\eta^{2}|\ln\varepsilon|}\right)m_{t}^{i,\varepsilon}(R)$
@@
As in the beginning of the proof of Proposition 4.8, we use the equation verified by $\omega^{\varepsilon}$ (3.2) and the decomposition of $u^{\varepsilon}$ (2.16), to state that $u_{L}^{\varepsilon}\cdot e_{r}=0$ implies
$\frac{\mathrm{d}}{\mathrm{d}t}\mu_{t}^{\varepsilon}(R,\eta)$ $=\frac{-1}{|\ln\varepsilon|}\int\big{(}1-W_{R,\eta}(x_{r}-r_{0}^{i,\varepsilon})\Big{)}\operatorname{div}\Big{(}(u^{\varepsilon}+F^{\varepsilon})(t,x)x_{r}\omega^{\varepsilon}(t,x)\Big{)}\mathrm{d}x$
$=\frac{-1}{|\ln\varepsilon|}\int W_{R,\eta}^{\prime}(x_{r}-r^{\varepsilon}_{0})e_{r}\cdot(u_{K}^{\varepsilon}+u_{R}^{\varepsilon}+F^{\varepsilon})(t,x)x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x=A_{1}+A_{2}$
@@
We start with $A_{1}$. Recalling from Lemmas 2.6 and 3.1 that $\|u_{R}^{\varepsilon}\|_{L^{\infty}}\leqslant C$ and $\|F^{\varepsilon}\|_{L^{\infty}}\leqslant C\sqrt{|\ln\varepsilon|}$, and that $W_{R,\eta}^{\prime}=0$ out of $[R,R+\eta]$ we have that
@@
We now compute $A_{2}$. First, we expand $u_{K}^{\varepsilon}$ and get that
@@
then we symmetrize by exchanging $x$ and $y$ to get that
@@
$|A_{22}+A_{23}|\leqslant\frac{C}{\eta^{2}|\ln\varepsilon|}\iint_{\mathcal{U}_{R}^{c}\times\mathcal{U}_{R/2}^{c}}x_{r}\omega^{\varepsilon}(t,x)y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}x\mathrm{d}y\leqslant\frac{C}{\eta^{2}|\ln\varepsilon|}m_{t}^{\varepsilon}(R)m_{t}^{\varepsilon}(R/2),$
@@
We now use the previous lemma to estimate the quantity $m_{t}^{i,\varepsilon}(R_{t}^{i,\varepsilon}/2)$. It is not possible to reach an estimate of order $\varepsilon^{\ell}$ in one step, so we use a double iterative method inspired by [11].
Starting from the weak concentration proved in Section 4, we get a weak concentration estimate of order $|\ln\varepsilon|^{-\ell}$.
@@
For every $\ell>0$, there exists a $T_{0}$ depending on $\ell$, that we denote by $T_{0,\ell}>0$, and $\varepsilon_{\ell}>0$, such that for every $\varepsilon\in(0,\varepsilon_{\ell}]$, $i\in\{1,\ldots,N\}$ and $t\in[0,T_{\varepsilon}]$, we have that
@@
In the statement of this lemma, we read the definition of $T_{\varepsilon}$ (3.6) as $T_{\varepsilon}\leqslant T_{0,\ell}$.
@@
by choosing $T_{0}$ small enough [^4] This assumption may seem not very natural. A possibility to remove it is to confine the solution in a strip centered not on $r_{0}^{\varepsilon}$, which is fixed, but on $X_{i,r}^{\varepsilon}(t)$, or even on $b_{i,r}^{\varepsilon}(t)$. However in the present paper, as other limitations exist on $T_{0}$, it does not appear have any real interest, while adding more technical difficulties.. From Lemma 5.3 and up to choosing $\varepsilon$ small enough, we have
@@
provided $R\in\left[\frac{d_{0}}{128\sqrt{|\ln\varepsilon|}},\frac{d_{0}}{64\sqrt{|\ln\varepsilon|}}\right]$ and $\varepsilon_{0}$ are small enough. Corollary 5.5 then implies that
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\mu_{t}^{\varepsilon}(R,\eta)\leqslant C\left(\frac{1}{\eta\sqrt{|\ln\varepsilon|}}+\frac{\ln|\ln\varepsilon|}{\eta^{2}|\ln\varepsilon|^{2}}\right)m_{t}^{\varepsilon}(R),$
for any $R\in\left[\frac{d_{0}}{128\sqrt{|\ln\varepsilon|}},\frac{d_{0}}{64\sqrt{|\ln\varepsilon|}}\right]$. Now we set $R_{0}=\frac{d_{0}}{64\sqrt{|\ln\varepsilon|}}$, $n=[\ln|\ln\varepsilon|]$ and $\eta=\frac{R_{0}}{2n}$ and we obtain that
@@
$\mu_{t}^{\varepsilon}(R,\eta)\leqslant\mu_{0}^{\varepsilon}(R,\eta)+C\,n\int_{0}^{t}m_{s}^{\varepsilon}(R)\mathrm{d}s$
@@
$\mu_{t}^{\varepsilon}(R,\eta)\leqslant\mu_{0}^{\varepsilon}(R,\eta)+Cn\int_{0}^{t}\mu_{s}^{\varepsilon}(R-\eta,\eta)\mathrm{d}s,$
@@
$\mu_{t}^{\varepsilon}(R_{0}-\eta,\eta)$ $\leqslant\mu_{0}^{\varepsilon}(R_{0}-\eta,\eta)+Cn\int_{0}^{t}\mu_{s_{1}}^{\varepsilon}(R_{0}-2\eta,\eta)\mathrm{d}s_{1}$
$\leqslant\mu_{0}^{\varepsilon}(R_{0}-\eta,\eta)+Cnt\mu_{0}^{\varepsilon}(R_{0}-2\eta,\eta)+(Cn)^{2}\int_{0}^{t}\int_{0}^{s_{1}}\mu_{s_{2}}^{\varepsilon}(R_{0}-3\eta,\eta)\mathrm{d}s_{2}\mathrm{d}s_{1}$
and iterating $n-1$ times eventually leads to
$\mu_{t}^{\varepsilon}(R_{0}-\eta,\eta)\leqslant\sum_{j=1}^{n-1}\mu_{0}^{\varepsilon}(R_{0}-j\eta,\eta)(Cnt)^{j-1}+(Cn)^{n-1}\int_{0}^{t}\ldots\int_{0}^{s_{n-2}}\mu_{s_{n-1}}^{\varepsilon}(R_{0}/2,\eta)\mathrm{d}s_{n-1}\ldots\mathrm{d}s_{1}$
and therefore, since $\mu_{0}^{\varepsilon}(R_{0}/2)=0$ (see the initial assumption (1.5)) and $\mu_{s}^{\varepsilon}\leqslant\gamma$ for every $s$, we obtain that
$\mu_{t}^{\varepsilon}(R_{0}-\eta,\eta)\leqslant\frac{(Cnt)^{n-1}}{(n-1)!}.$
@@
$|\ln\varepsilon|^{\ell}\mu_{t}^{\varepsilon}(R_{0}-\eta,\eta)$ $\leqslant\frac{1}{\sqrt{\pi(n-1)}}\exp\Big{(}(n-1)\big{(}\ln(Cnt)-\ln(n-1)+1\big{)}+\ell\ln|\ln\varepsilon|\Big{)}$
$\leqslant\frac{2}{\sqrt{\pi\ln|\ln\varepsilon|}}\exp\Big{(}\ln|\ln\varepsilon|\big{(}\ln(Ct)+2+\ell\big{)}\Big{)}.$
We see that if $T_{0,\ell}\leqslant\frac{1}{Ce^{\ell+2}}$, then this expression goes to 0 as $\varepsilon\to 0$ uniformly in $t\in[0,T_{\varepsilon}]$. By (6.1), this ends the proof. ∎
We iterate this procedure, starting now from the weak concentration proved in Lemma 6.4, to get a weak concentration estimate of order $\varepsilon^{\ell}$.
@@
By taking this time $R_{0}=\frac{d_{0}}{16\sqrt{|\ln\varepsilon|}}$, $\eta=\frac{R_{0}}{2n}$ and $n=[|\ln\varepsilon|]$, we get again that
@@
$\mu_{t}^{\varepsilon}(R_{0}-\eta,\eta)\leqslant\frac{(Cnt)^{n-1}}{(n-1)!}.$
@@
$\varepsilon^{-\ell}\mu_{t}^{\varepsilon}(R_{0}-\eta,\eta)$ $\leqslant\frac{1}{\sqrt{\pi(n-1)}}\exp\Big{(}(n-1)\big{(}\ln(Cnt)-\ln(n-1)+1\big{)}+\ell|\ln\varepsilon|\Big{)}$
$\leqslant\frac{2}{\sqrt{\pi|\ln\varepsilon|}}\exp\Big{(}|\ln\varepsilon|\big{(}\ln(Ct)+2+\ell\big{)}\Big{)}.$
which again goes to 0 if $T_{0,\ell}\leqslant\frac{1}{Ce^{\ell+2}}$. ∎
We are now able to use Lemma 6.2 to conclude the radial part of Proposition 6.1.
@@
Indeed, if it isn’t the case, by letting $t_{2}\in[t_{0},t_{1}]$ be the first time when $R_{t_{2}}^{\varepsilon}=f(t_{2})$, then there exists a point $x^{\varepsilon,0}\in\operatorname{supp}\omega^{\varepsilon}(t_{2})$ such that $|x^{\varepsilon,0}-r^{\varepsilon}_{0}|=R_{t_{2}}^{\varepsilon}$. However Lemma 6.2 is then in contradiction [^5] For more details on this contradiction argument, see for instance [22, Section 4.4]. with the fact that $t_{2}$ is the first such time, due to the fact that $f(t_{0})>R_{t_{0}}^{\varepsilon}$ and